# Nickel and Diming Your GAN: A Dual-Method Approach to Enhancing GAN Efficiency via Knowledge Distillation

### v1

@@ Abstract
In this paper, we address the challenge of compressing generative adversarial networks (GANs) for deployment in resource-constrained environments by proposing two novel methodologies: Distribution Matching for Efficient compression (DiME) and Network Interactive Compression via Knowledge Exchange and Learning (NICKEL). DiME employs foundation models as embedding kernels for efficient distribution matching, leveraging maximum mean discrepancy to facilitate effective knowledge distillation. Simultaneously, NICKEL employs an interactive compression method that enhances the communication between the student generator and discriminator, achieving a balanced and stable compression process. Our comprehensive evaluation on the StyleGAN2 architecture with the FFHQ dataset shows the effectiveness of our approach, with NICKEL & DiME achieving FID scores of 10.45 and 15.93 at compression rates of 95.73% and 98.92%, respectively. Remarkably, our methods sustain generative quality even at an extreme compression rate of 99.69%, surpassing the previous state-of-the-art performance by a large margin. These findings not only demonstrate our methodologies’ capacity to significantly lower GANs’ computational demands but also pave the way for deploying high-quality GAN models in settings with limited resources. Our code will be released soon.
@@ 1 Introduction
In addition to DiME, to exploit the characteristic of GAN that consists of a generator and a discriminator, we propose Network Interactive Compression via Knowledge Exchange and Learning (NICKEL). In GAN training, Lee et al. [29] has shown that the discriminator can provide more meaningful signals as feedback by learning the semantic knowledge of the generator. Inspired by Lee et al. , we not only distill knowledge directly between the generators (i.e.., DiME), but also distill knowledge from the more informative $G^{T}$ to the student discriminator ($D^{S}$) by transmitting knowledge between generators via the discriminator indirectly. By utilizing $G^{T}$, we obtain two distinct advantages, Firstly, from the onset of training, $D^{S}$ learns the rich semantic knowledge embedded within the $G^{T}$. Secondly, the $G^{T}$ provides a wealth of knowledge surpassing that of $G^{S}$. Furthermore, we observe that NICKEL enhances the stability of GAN compression (see Fig. 3). To the best of our knowledge, NICKEL is the first method that distills the knowledge from $G^{T}$ to $G^{S}$ via the feedback of $D^{S}$ for model compression.
Our experimental results show that DiME outperforms existing state-of-the-art compression methods through knowledge distillation between $G^{T}$ and $G^{S}$. By applying DiME to StyleGAN2, which has a baseline FID of 4.02, resulted in FID scores of 11.25 and 18.32 at compression rates of 95.87% and 98.92%, respectively. This compares favorably to the state-of-the-art method [16], which achieves FID scores of 14.01 and 22.23 at the same compression rate. This demonstrates the power of using foundation models as embedding kernels for knowledge distillation. In addition, by using NICKEL with DiME, we further enhance the FID scores to 10.45 and 15.93 with improved stability, setting a new standard in GAN compression performance. It is worth to note that we achieve a reasonable performance with the FID of 29.38 at the extreme compression rate of 99.69%, surpassing the previous state-of-the-art performance by a significant margin.
Our contributions can be summarized as follows:
  • With NICKEL & DiME, our final model further raises the bar of the state-of-the-art. We find that our method shows a stable convergence with competitive generative performance, even at the extremely high compression rates of 99.69% (Sec. 4).
@@ 2.1 GAN Compression
Recently, various GAN compression methods [31, 55, 13, 35, 32, 57, 16, 60, 58, 14] have been explored because directly applying compression methods, which were originally proposed for discriminative tasks, to GANs has led to significant performance degradation. However, most GAN compression methods have been explored in conditional GAN settings [31, 32, 60, 14], which are unsuitable for distribution matching in unconditional GAN compression (see Fig. 1). Occasionally, to address this problem, GAN compression methods have been explored [35, 16, 57], proposing better embedding spaces or distance metrics between $G^{T}$ and $G^{S}$. Wang et al. [55] proposed the GAN slimming (GS), a unified optimization framework, and emphasized that naïve application of compression methods leads to significant performance degradation due to the notorious instability of GANs. Liu et al. [35] proposed the content-aware GAN compression (CAGC) method, which focuses on only the contents of interest (e.g., object, face) to distill the knowledge, but this method requires additional costs due to manual labeling of contents. Li et al. [32] proposed the generator-discriminator cooperative compression (GCC) to maintain the nash-equilibrium between $G^{S}$ and $D^{S}$, but nash-equilibrium still cannot be maintained, in complex settings. Kang et al. [16] proposed the information-theoretic GAN compression (ITGC) by maximizing the mutual information between $G^{T}$ and $G^{S}$. ITGC requires a lot of computational costs due to the energy-based model and MCMC sampling. Xu et al. [57] proposed the StyleKD that focuses on the mapping network to achieve consistent outputs between $G^{T}$ and $G^{S}$. However, StyleKD can only be applied to networks based on StyleGAN. To address these issues, we propose Distribution Matching for Efficient compression (DiME), which matches the distribution between $G^{T}$ and $G^{S}$ via foundation kernels.
2.2 Foundation models
The foundation models [44, 45, 47, 41, 42, 1, 43, 2, 37] such as GPT [41, 42, 1], CLIP [43], and DINO [2, 37] are large-scale pre-trained models and are getting a lot of attention recently. Foundation models, which acquire general knowledge through large-scale training, are driving remarkable advancements on downstream tasks [61, 24, 34, 26]. For example, CLIP is a contrastive vision-language model trained so that positive pairs are clustered while negative pairs are scattered on a large dataset of 400 million image-text pairs. DINO is trained through self-supervised learning with self-distillation, passing all crops for the student and only the global view for the teacher to encourage local-to-global correspondences. In this paper, we utilize the CLIP and DINO, which have an image encoder, as kernels for better embedding spaces.
@@ 3 Method
This section introduces our two novel methodologies, DiME and NICKEL. Fig. 2 shows the overview schematic of our approach.
@@ 3.2 Network Interactive Compression via Knowledge Exchange and Learning
GAN utilizes a discriminator, which is a learnable network as the loss during the training of the generator. The performance of the generator is heavily influenced by the quality of feedback provided by the discriminator. GGDR [29] showed that during GAN training, the discriminator can learn semantic knowledge from the generator. Subsequently, the discriminator provides better feedback to the generator, thus improving the performance of the generator.
Inspired by GGDR, we propose NICKEL, which distills knowledge from $G^{T}$ into $D^{S}$ to provide more powerful feedback to $G^{S}$. NICKEL has advantages over simply applying GGDR to $G^{S}$ for two reasons. First, GGDR may struggle to provide meaningful information when $G^{S}$ resembles a random network during early training, whereas NICKEL can distill rich information from $G^{T}$ from the outset. Second, in GAN compression, due to the smaller network structure of $G^{S}$, GGDR cannot provide knowledge as rich as $G^{T}$. Therefore, we propose fine-tuning $D^{S}$ via NICKEL to learn information from $G^{T}$. However, fine-tuning $D^{S}$ using the NICKEL loss alone is insufficient to fully leverage the information from the pre-trained discriminator. Therefore, for adversarial learning, both $D^{T}$ and $D^{S}$ are employed.
The loss function of NICKEL can be formulated as follows:
@@ 4.1 Setups
4.1.1 Implementations.
We use the 20, 15, and 10 for $\lambda_{dino}$, $\lambda_{clip}$, and $\lambda_{\texttt{NICKEL}}$, respectively. We utilize the pretrained weights for CLIP [^1] https://openaipublic.azureedge.net/clip/models/afeb0e10f9e5a86da6080e35cf09123aca3b358a0c3e3b6c78a7b63bc04b6762/RN50.pt and DINO [^2] https://github.com/OpenGVLab/CaFo?tab=readme-ov-file as kernels. We use the same pruned generator as CAGC. We match 64 by 64 channels in StyleGAN2 and 8 by 8 channels in SNGAN for NICKEL on LH, HL, and HH components via Haar wavelet. To obtain the global features, we conduct 20,000 model inferences with the batchsize of 256.
Table 1: Comparison of FID scores of our methods and state-of-the-art compression methods on SNGAN for CIFAR-10. We compare the performance of various compression methods on SNGAN at compression rates of 74.88% and 90.85%. Distribution Matching for Efficient compression (DiME), in Eq. 1, shows superior performance compared to other state-of-the-art compression methods. Furthermore, NICKEL & DiME, which combines NICEKL (Eq. 2) and DiME, shows significant performance improvements. Particularly, at a compression rate of 74.88%, our method shows performance comparable to the full model.
@@
SNGAN CIFAR-10 Full model [36] 3.36B - 17.71
CAGC [35] 0.85B 74.88% 40.45
ITGC [16] 43.66
DiME 31.98
NICKEL & DiME 23.11
CAGC [35] 0.31B 90.85% 51.93
ITGC [16] 59.99
DiME 36.89
NICKEL & DiME 28.27
Table 2: Comparison of FID scores of our methods and state-of-the-art compression methods on StyleGAN2 for FFHQ and LSUN-CAT. We compare the performance of various compression methods on StyleGAN2 for FFHQ and LSUN-CAT datasets at various compression rates. Distribution Matching for Efficient compression (DiME), in Eq. 1, outperforms state-of-the-art GAN compression methods and shows a significant performance gap compared to state-of-the-art GAN compression methods at high compression rates. Our final method, combining NICKEL (Eq. 2) with DiME, significantly enhances the performance. Even NICKEL & DiME, which compresses the model by 98.92%, does not suffer significant performance degradation compared to the model compressed by 95.87% using ITGC. In other words, NICKEL & DiME is 3.83 times more efficient. Additionally, NICKEL & DiME incurs only a 9.68% performance decrease compared to the full model at a compression rate of 74.96%.
@@
StyleGAN2 FFHQ Full model [21] 14.90B - 4.02
ITGC [16] 3.73B 74.96% 5.27
@@
CAGC 0.61B 95.87% 14.29
ITGC 14.01
DiME 11.25
NICKEL & DiME 10.45
CAGC 0.16B 98.92% 23.05
@@
LSUN-CAT Full model [21] 14.90B - 8.19
GS 1.38B 90.73% 17.11
@@
4.1.2 Datasets.
We use FFHQ [20] and LSUN-CAT [59] datasets on StyleGAN2 [22], and CIFAR-10 [25] on SNGAN [36]. Resolution of FFHQ and LSUN-CAT is set to 256$\times$256.
@@
We follow the architecture and training setups of StyleGAN2 [21], except for augmentation (not used). Our pruned generators are identical to CAGC [35]. The architecture and training setups for SNGAN follow Kang et al. [17]. We use the official code of StyleGAN2-ADA-PyTorch [^3] https://github.com/NVlabs/stylegan2-ada-pytorch and StudioGAN [^4] https://github.com/POSTECH-CVLab/PyTorch-StudioGAN. For clear comparison, we reproduce state-of-the-art GAN compression methods based on the StyleGAN2-ADA-PyTroch official code.
4.1.4 Evaluation metrics.
We use the Fréchet Inception Distance (FID) [11] metric to evaluate the performance of the GANs, measuring both the quality and diversity of the generated images. FID compares the features extracted by Inception V3555https://nvlabs-fi-cdn.nvidia.com/stylegan2-ada-pytorch/pretrained/metrics/inception-2015-12-05.pt between real and generated images by assuming Gaussian distribution. Precision & Recall are metrics for calculating the fidelity and diversity of data by computing the manifold of the data through K-Nearest Neighbors. We use VGG16666https://nvlabs-fi-cdn.nvidia.com/stylegan2-ada-pytorch/pretrained/metrics/vgg16.pt to compute Precision & Recall. In this paper, we use 50,000 samples and set the neighborhood size to 3. We note the average of the five best FID scores.
@@ 4.2 Results
Additionally, to investigate the effectiveness of distillation considering the characteristics of GANs (i.e., NICKEL) beyond direct knowledge distillation (i.e., DiME), we combine NICKEL, as described in Eq. 2, and DiME. Tab. 1 and Tab. 2 show that NICKEL & DiME further improves the performance over that of DiME. In Tab. 2, our method achieves the FID score of 15.93 by compressing StyleGAN2 93-fold. This compares to ITGC, which attains the FID score of 14.01 with a 24-fold compression.
@@
StyleGAN2 FFHQ Full model [21] 14.90B - 4.02
GS [55] 0.05B 99.69% 184.33
@@ 4.4 Ablation Study
In this section, we study the power of kernel embedding spaces and the effect of components of our method.
CLIP and DINO embeddings. In Tab. 4, the CLIP embedding w/o global indicates using only CLIP as the embedding kernel for knowledge distillation. We observe significant challenges in achieving stable knowledge distillation when using only CLIP. On the other hand, when using only DINO as the embedding kernel, DINO embedding w/o global, we observe stable convergence and achieve the FID of 20.75. In addition, we observe that although the CLIP embedding space may pose challenges in achieving stable knowledge distillation, combining it with the DINO embedding space could lead to slight performance improvements.
@@
Fortunately, unlike $G^{S}$, the distribution of $G^{T}$ is fixed. Therefore, by precomputing the statistics— referred to as global features —through infinite sampling, we can achieve an infinitesimal sampling error $\epsilon_{teacher}$. As shown in Tab. 4, we find that utilizing global features leads to performance enhancement. In fact, this resembles the MMD critic, which is a stable metric for learning the distribution. Santos et al. [48] and Yeo et al. [58] noted that pretrained neural networks can be considered as characteristic kernels, and reducing the discrepancy of the mean between extracted features can be seen as the MMD critic. In this vein, DiME can be considered to stably match distributions between two generators. In summary, we observe that combining NICKEL & DiME shows the best performance.

### v2

@@ Abstract
In this paper, we address the challenge of compressing generative adversarial networks (GANs) for deployment in resource-constrained environments by proposing two novel methods: Distribution Matching for Efficient compression (DiME) and Network Interactive Compression via Knowledge Exchange and Learning (NICKEL). DiME employs foundation models as embedding kernels for efficient distribution matching, leveraging maximum mean discrepancy to facilitate effective knowledge distillation. NICKEL employs an interactive compression method that enhances the communication between the student generator and discriminator, achieving a balanced and stable compression process. Our comprehensive evaluation on the StyleGAN2 architecture with the FFHQ dataset shows the effectiveness of our approach, with NICKEL & DiME achieving FID scores of 10.45 and 15.93 at compression rates of 95.73% and 98.92%, respectively. Remarkably, our methods sustain generative quality even at an extreme compression rate of 99.69%, surpassing the previous state-of-the-art performance by a large margin. These findings not only show our methodologies’ capacity to significantly lower GANs’ computational demands but also pave the way for deploying high-quality GAN models in settings with limited resources. Our code is available at Nickel & Dime.
Keywords:
Model compression Generative models Compact models
@@ 1 Introduction
In addition to DiME, to exploit the characteristic of GAN that consists of a generator and a discriminator, we propose Network Interactive Compression via Knowledge Exchange and Learning (NICKEL). In GAN training, Lee et al. [33] has shown that the discriminator can provide more meaningful signals as feedback by learning the semantic knowledge of the generator. Inspired by Lee et al., we not only distill knowledge directly between the generators (i.e., DiME), but also distill knowledge from the more informative $G^{T}$ to the student discriminator ($D^{S}$) by transmitting knowledge between generators via the discriminator indirectly. By utilizing $G^{T}$, we obtain two distinct advantages, Firstly, from the onset of training, $D^{S}$ learns the rich semantic knowledge embedded within the $G^{T}$. Secondly, the $G^{T}$ provides a wealth of knowledge surpassing that of $G^{S}$. Furthermore, we observe that NICKEL enhances the stability of GAN compression (see Fig. 3). To the best of our knowledge, NICKEL is the first method that distills the knowledge from $G^{T}$ to $G^{S}$ via the feedback of $D^{S}$ for model compression.
Our experimental results show that DiME outperforms existing state-of-the-art compression methods through knowledge distillation between $G^{T}$ and $G^{S}$. By applying DiME to StyleGAN2, which has a baseline FID of 4.02, resulted in FID scores of 11.25 and 18.32 at compression rates of 95.87% and 98.92%, respectively. This compares favorably to the state-of-the-art method [21], which achieves FID scores of 14.01 and 22.23 at the same compression rate. This demonstrates the power of using foundation models as embedding kernels for knowledge distillation. In addition, by using NICKEL with DiME, we further enhance the FID scores to 10.45 and 15.93 with improved stability, setting a new standard in GAN compression performance. It is worth to note that we achieve a reasonable performance with the FID of 29.38 at the extreme compression rate of 99.69%, surpassing the previous state-of-the-art performance by a significant margin. Our contributions can be summarized as follows:
@@
  • With NICKEL & DiME, our final model further raises the bar of the state-of-the-art. Our method shows a stable convergence with competitive performance, even at the extremely high compression rates of 99.69% (Sec. 4).
  • Last but not least, we standardize and benchmark GAN compression methods using official codes, ensuring future compatibility and reproducibility.
@@ 2.1 GAN Compression
Most GAN compression methods have been explored in conditional GAN settings [35, 36, 62, 17], which are unsuitable for distribution matching in unconditional GAN compression (see Fig. 1). Occasionally, to address this problem, GAN compression methods have been explored [39, 21, 59], proposing better embedding spaces or distance metrics between $G^{T}$ and $G^{S}$. Wang et al. [57] proposed the GAN slimming (GS), a unified optimization framework, and emphasized that naïve application of compression methods leads to significant performance degradation due to the notorious instability of GANs. Liu et al. [39] proposed the content-aware GAN compression (CAGC) method, which focuses on only the contents of interest (e.g., object, face) to distill the knowledge, but this method requires additional costs due to manual labeling of contents. Li et al. [36] proposed the generator-discriminator cooperative compression (GCC) to maintain the nash-equilibrium between $G^{S}$ and $D^{S}$, but nash-equilibrium still cannot be maintained, in complex settings. Kang et al. [21] proposed the information-theoretic GAN compression (ITGC) by maximizing the mutual information between $G^{T}$ and $G^{S}$. ITGC requires a lot of computational costs due to the energy-based model and MCMC sampling. Xu et al. [59] proposed the StyleKD that focuses on the mapping network to achieve consistent outputs between $G^{T}$ and $G^{S}$. However, StyleKD can only be applied to networks based on StyleGAN. To address these issues, we propose Distribution Matching for Efficient compression (DiME), which matches the distribution between $G^{T}$ and $G^{S}$ via foundation kernels.
@@ 3.2 Network Interactive Compression via Knowledge Exchange and Learning
GAN utilizes a discriminator, which is a learnable network as the loss during the training of the generator. The performance of the generator is heavily influenced by the quality of feedback provided by the discriminator. GGDR [33] showed that during GAN training, the discriminator can learn semantic knowledge from the generator. Subsequently, the discriminator provides better feedback to the generator, thus improving the performance of the generator. Inspired by GGDR, we propose NICKEL, which distills knowledge from $G^{T}$ into $D^{S}$ to provide more powerful feedback to $G^{S}$. NICKEL has advantages over simply applying GGDR to $G^{S}$ for two reasons. First, GGDR may struggle to provide meaningful information when $G^{S}$ resembles a random network during early training, whereas NICKEL can distill rich information from $G^{T}$ from the outset. Second, in GAN compression, due to the smaller network structure of $G^{S}$, GGDR cannot provide knowledge as rich as $G^{T}$. Therefore, we propose fine-tuning $D^{S}$ via NICKEL to learn information from $G^{T}$. However, fine-tuning $D^{S}$ using the NICKEL loss alone is insufficient to fully leverage the information from the pre-trained discriminator. Therefore, for adversarial learning, both $D^{T}$ and $D^{S}$ are employed. The loss function of NICKEL can be formulated as follows:
@@ 4.1 Setups
4.1.1 Implementation details.
We set 20, 15, and 10 for $\lambda_{dino}$, $\lambda_{clip}$, and $\lambda_{\texttt{NICKEL}}$, respectively. We use the same pruned generator as CAGC, and for CLIP and DINO, we use the pretrained weights. [^1] CLIP.pt and DINO.pt For NICKEL, we match feature maps in 1/4 resolution channels of the generator’s output (e.g., 64x64 feature maps for 256 resolution) on LH, HL, and HH components of Haar wavelet. To obtain the global features, we conduct 20,000 model inferences with the batchsize of 256.
Table 1: Comparison of FID scores of our methods (DiME and NICKEL & DiME) and state-of-the-art compression methods on StyleGAN2 for various datasets. The dagger symbol indicates the performance of the official model provided on the original paper’s GitHub repository. The right side of the arrow shows the results obtained using NVIDIA’s official FID measurement code. Our reproduced models achieve higher performance compared to the official models.
@@
StyleGAN2 FFHQ (256$\times$256) (1) Full model [23] 24.77M 14.90B - 4.02
(2) Full model† 30.03M 45.12B - 4.5
@@
[…]
CAGC 8.72M 3.73B 74.96% 5.24
ITGC [21] 5.27
DiME 5.00
@@
CAGC 2.69M 0.16B 98.92% 23.05
@@
FFHQ (1024$\times$1024) Full model 30.37M 74.27B - 2.74
Full model† 49.1M 74.3B - 2.7
[…]
NICKEL & DiME 5.65M 6.99B 90.59% 6.41
LSUN Church (256$\times$256) Full model 30.03M 45.12B - 3.97
CAGC 5.57M 4.12B 90.87% 4.50
StyleKD 4.47
[…]
NICKEL & DiME 3.94
DDPM LSUN Church (256$\times$256) Full model† 113.7M 497.4B - 10.6
SPDM (t=100)†[7] 63.2M 277.6B 44.19% 13.9
Table 2: Comparison of FID scores of our methods (DiME and NICKEL & DiME) and state-of-the-art compression methods on various architectures for various datasets.
[…]
SNGAN CIFAR-10 (32$\times$32) Full model [41] 4.28M 3.36B - 17.71
CAGC [39] 1.20M 0.85B 74.88% 40.45
ITGC [21] 43.66
DiME 31.98
NICKEL & DiME 23.11
CAGC 0.50M 0.31B 90.85% 51.93
ITGC 59.99
DiME 36.89
NICKEL & DiME 28.27
BigGAN CIFAR-10 (32$\times$32) Full model [1] 8.83M 3.83B - 10.66
CAGC 0.89M 0.37B 90.45% 26.32
ITGC 27.78
NICKEL & DiME 26.66
StyleGAN2 CelebA (128$\times$128) Full model [23] 24.53M 11.27B - 2.70
CAGC 3.08M 0.27B 97.60% 10.05
ITGC 10.09
StyleKD 10.41
NICKEL & DiME 7.56
AFHQ (512$\times$512) Full model 30.28M 59.67B - 3.01
NICKEL & DiME 10.30M 14.94B 74.96% 3.17
LSUN CAT (256$\times$256) Full model 24.77M 14.90B - 8.19
GS 4.96M 1.38B 90.73% 17.11
@@
4.1.2 Datasets and Evaluation Metrics.
We use CelebA [40], FFHQ [24], AFHQ [4], LSUN-Church and LSUN-CAT [61] datasets, and CIFAR-10 [28]. se Fréchet Inception Distance (FID) [14] and Precision & Recall [30, 42, 27] to evaluate the performance of the GANs, measuring both the quality and diversity of the generated images. We note the average of the five best FID scores. , we use the fvcore library to measure FLOPs, ensuring consistency with reported FLOPs and enabling fair comparisons with existing works. [^2] If you noticed any discrepancy in FLOPs, the discrepancy stems from (a) the typo in the original CAGC paper, where FLOPs were reported as 4.1M instead of 4.1B—a mistake recognized in subsequent papers; and (b) different libraries used. The reported FLOPs of 4.1B at 90.87% compression can be reproduced via fvcore library. In contrast, when used torchprofile library with the official NVIDIA StyleGAN2 code, it gives the FLOPs of 5.33B at 90.73%.
@@
We follow the architecture and training setups of StyleGAN2 [23], except for augmentation (not used). Our pruned generators are identical to CAGC [39]. The architecture and training setups for BigGAN and SNGAN follow Kang et al. [19]. We use the official code of StyleGAN2-ADA-PyTorch and StudioGAN. We also compare Structural Pruning for Diffusion Models (SPDM) [7].
4.2 Benchmarking and Reproducibility in GAN Compression
The performance metrics for previous GAN compression techniques were often based on outdated and unofficial code. To address this and for future compatibility and reproducibility in the field, we reimplement the state-of-the-art GAN compression methods using NVIDIA’s official StyleGAN2-ADA code. The official code shows comparable performance to the unofficial code, but with more efficient FLOPs and fewer parameters (see (1) and (2) in Tab. 1). Additionally, we provide results re-evaluated using NVIDIA’s official FID code, utilizing trained weights from the original GitHub repository of the baselines (indicated by the right of the arrow of (3)-(7) in Tab. 1). While the re-evaluated FID scores for 1024$\times$1024 FFHQ and LSUN-Church datasets are similar to the reported FID, we find that the FID for the 256$\times$256 FFHQ dataset shows a significant performance gap (see (3) and (4) in Tab. 1). This discrepancy arises because the reported FID scores were not calculated by computing the feature embedding directly but by using the feature embedding provided by CAGC’s custom FID code. We hypothesize that the feature embedding provided by CAGC may be biased for the 256$\times$256 FFHQ dataset.
@@ 4.3 Results
Additionally, to investigate the effectiveness of distillation considering the characteristics of GANs (i.e., NICKEL) beyond direct knowledge distillation (i.e., DiME), we combine NICKEL, as described in Eq. 2, and DiME. Tab. 1 and Tab. 2 show that NICKEL & DiME further improves the performance over that of DiME. In Tab. 1, our method achieves the FID score of 15.93 by compressing StyleGAN2 93-fold. This compares to ITGC, which attains the FID score of 14.01 with a 24-fold compression. Furthermore, as shown in Tab. 1, our method not only obtains better computational efficiency but also shows superior performance compared to the state-of-the-art diffusion-model pruning method for the LSUN-Church dataset. These results indicate the continued significance of GAN compression research. As shown in Tab. 1 and Tab. 2, our method can be applied to various GAN architectures and show similar trends for various datasets.
@@
StyleGAN2 FFHQ (256$\times$256) Full model [23] 24.77M 14.90B - 4.02
GS [57] 2.35M 0.05B 99.69% 184.33
@@ 4.5 Ablation Study
CLIP and DINO Embeddings. In Tab. 4, the CLIP embedding w/o global indicates using only CLIP as the embedding kernel for knowledge distillation. We observe significant challenges in achieving stable knowledge distillation when using only CLIP. In contrast, when using only DINO as the embedding kernel, DINO embedding w/o global, we observe stable convergence and achieve the FID of 20.75. In addition, we observe that although the CLIP embedding space may pose challenges in achieving stable knowledge distillation, combining it with the DINO embedding space could lead to slight performance improvements.
@@
Fortunately, unlike $G^{S}$, the distribution of $G^{T}$ is fixed. Therefore, by precomputing the statistics— referred to as global features —through infinite sampling, we can achieve an infinitesimal sampling error $\epsilon_{teacher}$. As shown in Tab. 4, we find that utilizing global features leads to performance enhancement. In fact, this resembles the MMD critic, which is a stable metric for learning the distribution. Santos et al. [50] and Yeo et al. [60] noted that pretrained neural networks can be considered as characteristic kernels, and reducing the discrepancy of the mean between extracted features can be seen as the MMD critic. In this vein, DiME can be considered to stably match distributions between two generators.